The sum of Carol's and Sandy's ages is 47. In 10 years Carol will be 10 years older than Sandy was 7 years ago. Find their present ages.
Answer by AnlytcPhil(1806)


The sum of Carol's and Sandy's ages is 47. In 10 years Carol will be
10 years older than Sandy was 7 years ago. Find their present ages.
`
The sum of Carol's and Sandy's ages (C+S) is 47. In 10 years Carol (C+10)
will be 10 years older than(10 more than) Sandy was 7 years ago (S-7).
`
(C+S) = 47
(C+10) = 10 more than (S-7), which means "(S-7) plus 10"
`
or
`
C + S = 47
C + 10 = (S - 7) + 10
`
Can you solve this system, after simplifying the second one?
`
Answer: Carol is 20 and Sandy is 27
`
Edwin